You have a credit card that charges an APR of 27%. You make a variable amount of charges each month (see below) but you make the same payment amount each month: $250. If you start with a balance of $0 and borrow the amounts indicated below, what would be the balance owed at the end this year.

Month Borrowed amount

June $280

July $320

August $400

September $300

October $200

November $180

December $350

January $150

February $250

March $210

April $200

May $280

Balance after these 12 months: ?
1
Expert's answer
2019-05-29T09:26:41-0400

Monthly rate: 0.27/12=0.0225

June: 0$- $280+$250=-$30; interest -$30*0.0225=-$0.68

July: (-$30-$0.68) -$320+$250=-$100.675; interest -$100.675*0.0225=-$2.27

August: (-$100.675-$2.27)-$400+$250=-$252.94; interest -$252.94*0.0225=-$5.69

September: (-$252.94-$5.69)-$300+$250=-$308.63; interest -$308.63*0.0225=-$6.94

October: (-$308.63-$6.94)-$200+$250=-$265.58; interest -$265.58*0.0225=-$5.98

November: (-$265.58-$5.98)-$180+$250=-$201.55; interest -$201.55 *0.0225=-$4.53

December:(-$201.55-$4.53)-$350+$250=-$306.09; interest -$306.09*0.0225=-$6.89

January: (-$306.09-$6.89)-$150+$250=-$212.97; interest -$212.97*0.0225=-$4.79

February: (-$212.97-$4.79)-$250+$250=-$217.76; interest -$217.76*0.0225=-$4.90

March: (-$217.76-$4.90)-210+$250=-$182.66; interest -$182.66*0.0225=-$4.11

April: (-$182.66-$4.11)-$200+$250=-$136.77; interest -$136.77*0.0225=-$3.08

May:(-$136.77-$3.08)- $280+$250=-$169.85; interest -$169.85*0.0225=-$3.82

Balance in the beginning of June:-$169.85-$3.82=-$173.67